Can someone Please Help me Solve!!
To determine the number of deer in a game preserve, a conservationist catches 569 deer, tags them and lets them loose. Later, 868 deer are caught; 217 of them are tagged. How many deer are in the preserve?

Answer by chessace(471)

One has to make quite a few dicey assumptions to answer this, but the only number one can compute is via:
569 is f fraction of total pop x.
f = 569 / x.
IF yadda yadda, 217 = same fraction of 868.
f = 217 / 868.
So 569 / x = 217 / 868
cross multiply etc to get x = 569 * 868 / 217.
x = 2276
Quick check: f is about 1/4 (actualy, f = 1/4).